Archie’s gas tank is
1
/
3
full
.
If Archie adds 3 gallons of gas to the tank, it will be
˝
full
.
What is thecapacity in gallons of Archie’s tank?

Answer by hamsanash1981@gmail.com(151)

Let the total gallons to fill the tank be x
when 3 gallons of gas added to 1/3 of x it becomes 1/2 of x
=> x/3+3 = x/2
=> (x+9)/3 = x/2
=> 2x+18 = 3x
=> x=18
The capacity of the tank in gallons is 18.